OpenStudy (anonymous):

Charla and David are observing the leaf growth rates of their plants. Charla’s plant starts on day 1 with 43 leaves and gets 1 new leaf a day. David’s plant starts on day 1 with 27 leaves and gets 5 leaves a day. How many days will it take David’s plant to have as many leaves as Charla’s plant? A. 7 B. 4 C. 9 D. 8

OpenStudy (anonymous):

4

OpenStudy (anonymous):

let x be number of days we have 43 + 1x = y 27 + 5x = y where y is day when they are equal subtract 4x - 16 = 0 x = 4